OpenStudy (anonymous):

the future of my high school career is dependent upon these questions... not really but ya know 1. 7^0 2. 2^-2 3. 2^-4•18^0 4. (2•5)^-2 5. 3^0•10^-4 6. a^-4 7. 3a^2b^-1 SAVE MEH

OpenStudy (experimentx):

1 1/4 = 0.25 1/16 = 4/25 = 1/10000 1/a^4 3a^2/b

OpenStudy (callisto):

First you need to know (i)a^0 = 1, and (ii) a^(-n) = 1/(a^n) then from (i) , you can solve 1. for 2, put a =2 and n=2 into a^(-n) = 1/(a^n) for 3, 2^-4•18^0 =[ 1/(2^4) ] (1) for 4, (2•5)^-2 = 10^-2 = 1/(10^2) = 1/100 for 5, 3^0•10^-4 = 1 (10^-4) = see experimentX 's answer for 6, put n = 4 into a^(-n) = 1/(a^n) for 7, 3a^2b^-1 = 3 (a^2) / b (same as experimentX 's answer) Well. experimentX saved you :P

OpenStudy (experimentx):

thanks for explanation

OpenStudy (callisto):

lol i seem like copying your answers ..

OpenStudy (experimentx):

still i you saved me ... i got complaints for not explaining properly.

OpenStudy (callisto):

Nah.. next time , save yourself, you can do it !

OpenStudy (experimentx):

yeah ... sure i hope so.. lol